### v4

# Large Language Model Sentinel: LLM Agent for Adversarial Purification

Abstract
Over the past two years, the use of large language models (LLMs) has advanced rapidly. While these LLMs offer considerable convenience, they also raise security concerns, as LLMs are vulnerable to adversarial attacks by the well-designed textual perturbations. In this paper, we introduce a novel defense technique named Large LAnguage MOdel Sentinel (LLAMOS), which is designed to enhance the adversarial robustness of LLMs by purifying the adversarial textual examples before feeding them into the target LLM. Our method comprises two main components: a) Agent instruction, which can simulate a new agent for adversarial defense, altering minimal characters to maintain the original meaning of the sentence while defending against attacks; b) Defense guidance, which provides strategies for modifying clean or adversarial examples to ensure effective defense and accurate outputs from the target LLMs. Remarkably, the defense agent demonstrates robust defensive capabilities even without learning from adversarial examples. Additionally, we conduct an intriguing adversarial experiment where we develop two agents, one for defense and one for attack, and engage them in mutual confrontation. During the adversarial interactions, neither agent completely beat the other. Extensive experiments on both open-source and closed-source LLMs demonstrate that our method effectively defends against adversarial attacks, thereby enhancing adversarial robustness.
Keywords:
Adversarial Robustness Adversarial Purification Large Language Model (LLM) LLM Agent.
1 Introduction
Figure 1: Illustration of adversarial attack and purification on LLMs. (a) The clean example $x$ is perturbed to the adversarial example $x^{\prime}$, causing misclassification. (b) LLAMOS purifies $x^{\prime}$ to the purified example $\hat{x}$, restoring the correct label.
Large language models (LLMs) have garnered significant attention due to their impressive performance across a wide range of natural language tasks [18]. The pre-trained LLMs, such as Meta’s LLAMA [31, 32] and OpenAI’s ChatGPT [21], have become essential foundations for AI applications in various sectors such as healthcare, education, and visual tasks [22, 7, 24]. Despite their widespread use and convenience, concerns about the security of these models are increasing. Specifically, LLMs have been shown to be vulnerable to adversarial textual examples [34, 41], which involve subtle modifications to textual content that maintain the same meaning for humans but completely change the prediction results to LLMs, often with severe consequences.
To achieve robust defense against adversarial attacks on LLMs, a prevalent strategy is fine-tuning the LLMs with adversarial examples to enhance model alignment [25, 37]. LLM-based adversarial fine-tuning (AFT) can be implemented either through in-context learning [4, 40] or by optimizing the parameters of pre-trained LLMs using adversarial examples [3, 11]. However, LLM-based AFT methods necessitate additional computational resources and training time. Achieving robust and reliable LLMs typically requires significant costs [6], which is prohibitive for ordinary users. Additionally, due to the discrete nature of textual information, these adversarial examples can have substitutes in any token of the sentence, with each having a large candidate list [9]. This leads to a combinatorial explosion, making the application of AFT methods challenging or resulting in poor generalization when trained on a limited dataset of adversarial examples. Consequently, developing an efficient and user-friendly robust LLM system remains a huge challenge and an urgent issue that continues to be addressed.
In this paper, focusing on adversarial textual attacks targeting LLM-based classification tasks, we propose a novel defense technique named Large LAnguage MOdel Sentinel (LLAMOS), which utilizes the LLM as a defense agent for adversarial purification, as illustrated in Figure 1a. To streamline our explanation, we condense certain details in Figure 1b, with comprehensive instructions provided in the Methods Section. Specifically, LLAMOS comprises two components: Agent instruction, which can simulate a new agent for adversarial defense, altering minimal characters to maintain the original meaning of the sentence, and Defense guidance, which provides strategies for modifying clean or adversarial examples to ensure effective defense and accurate outputs from the target LLMs. LLAMOS serves as a pre-processing method aiming to eliminate harmful information from potentially attacked textual inputs before feeding them into the target LLM for classification. In contrast to the AFT method, the LLM-based AP method functions as an additional module capable of defending against adversarial attacks without necessitating fine-tuning of the target LLM.
We comprehensively evaluate the performance of our method on GLUE datasets, conducting experiments with both representative open-source and closed-source LLMs, LLAMA-2 [32] and GPT-3.5 [21]. The experimental results demonstrate that the LLM-based AP method effectively defends against adversarial attacks. Specifically, our proposed method achieves a maximum reduction in the attack success rate (ASR) by up to 45.59% and 37.86% with LLAMA-2 and GPT-3.5, respectively. Additionally, we observe that the initial defense agent fails to achieve the expected results under some obvious attacks. Therefore, we employ the in-context learning [4] to further optimize the defense agent, significantly enhancing the defense capabilities almost without adding any additional costs. Finally, we conduct an intriguing online adversarial experiment, creating an adversarial system using two LLM-based agents (one for defense and one for attack) along with a target LLM for classification.
2 Related Work
2.0.1 Adversarial Attack
Deep neural networks (DNNs) are vulnerable to adversarial examples [30], which are generated by adding small, human-imperceptible perturbations to natural examples, but completely change the prediction results to DNNs [5, 13]. With the rapidly increasing applications of LLMs [22, 7, 24], security concerns have emerged as a critical area of research [9, 43], with researchers increasingly focusing on adversarial attacks targeting LLMs. In a similar setup to DNNs, for LLMs, attackers manipulate a small amount of text to change the output of the target LLM while maintaining the semantic information for humans [35, 41]. Presently, addressing the security issues surrounding LLMs is of paramount importance and requires urgent attention.
2.0.2 Adversarial Defense
There are two main defense techniques on traditional DNNs, including adversarial training (AT) [5] and adversarial purification (AP) [26]. Unlike traditional DNNs, retraining LLMs is nearly impossible due to cost issues [9]. Therefore, most methods enhance the robustness of LLMs through adversarial fine-tuning (AFT) [40, 11]. While AFT can effectively defend against attacks, it remains susceptible to unseen attacks whose adversarial examples that the LLMs have not previously learned [9]. Additionally, even with fine-tuning, training the LLMs will still consume a significant cost [6, 3]. Adversarial purification (AP) aims to purify adversarial examples before feeding them into the target model, which has emerged as a promising defense method [26, 13]. Compared with the AT or AFT method, the AP method utilizes an additional model that can defend against unseen attacks without retraining the target model [13, 9]. In some traditional computer vision and natural language processing tasks, researchers have started using LLMs for adversarial purification [27, 10, 19], but the security issues of LLMs themselves have not been deeply considered. Therefore, we propose a novel LLM defense technique named LLAMOS to purify the adversarial textual examples before feeding them into the target LLM, aiming to improve the robustness.
2.0.3 Large Language Model Agent
The LLM agent is a new research direction that has emerged in recent years [20, 17]. This novel type of agent is capable of interacting with humans in natural language, leading to a significant increase in applications across fields such as chatbots, natural sciences, robotics, and workflows [12, 15]. Furthermore, LLMs have demonstrated promising zero-shot/few-shot planning and reasoning capabilities across various configurations [29], covering specific environments and reasoning tasks [42]. In this paper, we introduce a new variant of the LLM agent designed specifically to purify adversarial textual examples generated by attacks.
3 Preliminary
3.0.1 Adversarial Attcks on LLMs
Given a target LLM $f_{t}$ with task instruction, input $x$, and correct label $y$, attackers aim to find the adversarial examples $x^{\prime}$ that can fool the target LLM $f_{t}$ on classification tasks. The adversarial examples $x^{\prime}$ can be obtained by the LLM itself $f_{atk}$ with different system prompt [41],
$x^{\prime}=f_{atk}(x)=x+\delta,\quad f_{t}(x^{\prime})=y^{\prime}\neq y,$
where $\delta$ represents textual perturbations from a series of candidate sets for modifications, which are made at the character level, word level, or sentence level. In a specific instance, the system prompt of $f_{t}$ can be: “Analyze the tone of this statement and respond with either ‘positive’ or ‘negative’.” and the system prompt of corresponding $f_{atk}$ can be: “Your task is to generate a new sentence that keeps the same semantic meaning as the original one but be classified as a different label.” There are more details in Appendix B.
3.0.2 Evaluations of LLMs Robustness
To evaluate the effectiveness of the defense method, we follow the setting from [36, 41], using the attack success rate (ASR) and traditional robust accuracy (RA) on the adversarial examples as measures of the robustness of the defense method.
$\text{ASR}=\frac{\sum_{(x,y)\in D}\mathbbm{1}\left[f_{t}(x^{\prime})\neq y\right]\cdot\mathbbm{1}\left[f_{t}(x)=y\right]}{\sum_{(x,y)\in D}\mathbbm{1}\left[f_{t}(x)=y\right]},\quad\text{RA}=\frac{1}{N}\sum\mathbbm{1}\left[f_{t}(x^{\prime})=y\right],$ (1)
where $\mathbbm{1}[\cdot]\in\{0,1\}$ is an indicator function. $D$ is the original test dataset and $D^{\prime}$ is the adversarial example dataset, and $N$ is the number of examples. The lower the ASR, the higher the RA, indicating greater model robustness.
4 Methods
We propose a novel defense technique for large language model-based adversarial purification (LLAMOS), which purifies adversarial textual examples by the LLM-based defense agent before feeding them into the target LLM. Firstly, we outline the overall pipeline of LLAMOS. Subsequently, we further augment the defense agent using in-context learning. Finally, we present the design of the adversarial system, incorporating the defense agent, attack agent, and target LLM.
4.1 Overview of LLAMOS
To defend against adversarial textual attacks targeting LLM-based classification tasks, we propose Large LAnguage MOdel Sentinel (LLAMOS) that employs the LLM as a defense agent for adversarial purification. LLAMOS comprises two components: Agent instruction and Defense guidance. Next, we introduce the overall pipeline in sequential order.
In this paper, we utilize the existing LLM denoted by target LLM $f_{t}$ as the classifier. Given an adversarial example $x^{\prime}$, the target LLM $f_{t}$ outputs the incorrect label $y^{\prime}=f_{t}(x^{\prime})$, while after purification, the purified example $\hat{x}=g_{llm}(x^{\prime})$ can be predicted as the correct label $y=f_{t}(g_{llm}(x^{\prime}))$. To achieve this, we design prompts for generating a defense agent $g_{llm}$ as described in the following.
{mdframed}
# Defense Agent Instruction
To begin, let me provide a brief overview of the input text: [Input Description]. The classification task for these sentences is [Task Description]. However, be aware that these sentences might be susceptible to adversarial attacks, which could lead to an incorrect label. Note that not all sentences will be affected by the attacks. Your task is to generate a new sentence that replaces the original one, which must satisfy the following conditions: [Defense Goal].
# Defense Guidance
You can complete the task using the following guidance: [Defense Guidance]. Input: [Input]. Now, let’s start the defense process and only output the generated sentence.
Input Description: The format of Input $x_{in}\in\{D_{i},i=1...6\}$ varies significantly across different datasets $D_{i}$, necessitating tailored input formats $x_{in}$ to correspond with the specific structure and content of each dataset, details in Table 10. For instance, the SST-2 dataset [28] typically consists of a single sentence per data point. On the other hand, the MNLI dataset [39] is structured to include pairs of sentences labeled as premise and hypothesis.
Task Description: Similar to the input description, the tasks associated with each dataset $D_{i}$ are distinct. As illustrated earlier, SST-2 focuses on determining the sentiment of a given sentence, making it a straightforward classification task. Conversely, MNLI presents a more complex task of natural language inference, where the relationship between a pair of sentences must be discerned and classified correctly. Detailed input and task descriptions are provided in Table 9.
Defense Goal: Based on traditional adversarial purification [26, 13, 14], we design the goal for LLAMOS as follows: “Keeping the semantic meaning of the new sentence the same as the original one; For natural examples, the new sentence should remain unchanged. For adversarial examples, modify the sentence so that it is classified as the correct label, effectively reversing the adversarial effect.”
Defense Guidance: The defense guidance offers specific instructions to the defense agent on how to modify the input text to ensure effective defense and accurate outputs from the target LLM. In designing our guidance, we considered attacks at various levels [41], including character, word, and sentence levels, which are presented in Table 1. These guidances are not rigidly fixed; they can be fine-tuned according to specific tasks.
After the defense agent generates the new sentence, the purified example $\hat{x}=g_{llm}(x^{\prime})$ is input into the target LLM $f_{t}$ for classification.
Figure 2: Illustration of adversarial system with attack and defense agents.
Table 1: The guidances for defense agent.
Idx. [Defense Guidance]
1 Modify as few characters as possible.
2 Correct any clear spelling errors.
3 Eliminate redundant symbols.
4 If necessary, feel free to replace, delete,
add words, or adjust the word order.
5 Improve structure for better readability.
6 Ensure sentence is coherent and logical.
4.2 Enhencing the Defense with In-Context Learning
In the initial defense agent, the defense guidance relies on common sense, which may result in poor performance against some special attacks, even when the attacker adds obvious characters. To address this limitation, we introduce in-context learning (ICL) [4] to further optimize the defense agent by,
{mdframed}
# In-Context Learning
The new sentence still contains a lot of harmful content caused by adversarial attacks, such as [Specific Guidance]. Please consider these contents and output a new sentence for me. Input: [Input]. Now, let’s start the defense process and only output the generated sentence.
The specific guidance is designed to assist the defense agent in better understanding an attack and generating a new sentence capable of effectively defending against the attack. These guidances can be fine-tuned to address specific attacks and can be incorporated into the defense agent as needed. Through in-context learning, the defense agent can be continuously optimized, significantly enhancing its performance almost without adding any additional costs.
4.3 Adversarial System with Multiple LLMs
In this section, we devise an adversarial system involving multiple LLMs. Given that our method introduces a defense agent against attackers, a natural idea is to then create an attack agent to counter the defender. The attack agent is tasked with generating adversarial examples from purified examples to deceive the target LLM once more. To accomplish this, we design prompts for generating an attack agent $g_{atk}$ by, {mdframed} # Attack Agent Instruction
To begin, let me provide a brief overview of the input text: [Input Description]. The classification task for these sentences is [Task Description]. Your task is to generate a new sentence that replaces the original one, which must satisfy the following conditions: [Attack Instruction].
# Attack Guidance
For example, the original sentence [Purified Example] is classified as [Correct Label]. You should generate a new sentence which is classified as [Incorrect Label]. Input: [Input]. Now, let’s start the attack process and only output the generated sentence.
The prompt structure of attack agent and defense agent is basically the same, although there are some differences in details. The input description of attack agent includes correct label $y$, and the input format is $(x_{in},y)\in D$. The attack instruction is “1. The new sentence should be classified as the opposite of the ‘correct label’. 2. Change at most two letters in the sentence.” Finally, we provide a specific example to help the attack agent better understand the attack task.
Then, we combine the defense agent and attack agent to build an adversarial system, as illustrated in Figure 2. In the adversarial system, the purified examples can be attacked again by the attack agent, and likewise, the adversarial examples can also be purified by the defense agent. They continuously counter each other, much like adversarial training [5].
5 Experiments
We conduct extensive experiments on GLUE datasets to evaluate the effectiveness of the proposed method. Specifically, our method reduces the attack success rate by up to 37.86% with GPT-3.5 and 45.59% with LLAMA-2, respectively.
5.1 Experimental Setup
5.1.1 Datasets
The experiments are conducted on six tasks in GLUE datasets [33], including SST-2 [28], RTE [2], QQP [38], QNLI [23], MNLI-mm, MNLI-m [39]. The detailed descriptions are provided in Appendix A.
5.1.2 Adversarial Attacks
We evaluate our method against PromptAttack [41], which is a powerful attack that combines nine different types of attacks, as illustrated in Table 11. Furthermore, [41] introduces the few-shot (FS) [16] and ensemble (EN) strategies [1] to boost the attack power, details in Appendix B.
5.1.3 Evaluation Metrics
We evaluate the performance of defense methods using two metrics: attack success rate (ASR) and robust accuracy (RA). A lower ASR or higher RA indicates greater model robustness.
5.1.4 Experiment Details
The experiments in this paper are conducted using ‘GPT-3.5-Turbo-0613’ and ‘LLAMA-2-7b’. For GPT-3.5, we purchase OpenAI’s API service and conduct testing experiments with the ‘openai’ package in Python. For LLAMA-2, we deploy it locally on NVIDIA RTX A6000 and utilize the available checkpoint published by MetaAI from HuggingFace.
Table 2: The attack success rate (ASR) defense against PromptAttack-EN and PromptAttack-FS-EN on the GLUE dataset with GPT-3.5.
Attacks LLAMOS SST-2 RTE QQP QNLI MNLI-mm MNLI-m Avg.
PA-EN $\times$ 56.00 34.30 37.03 40.39 43.51 44.00 42.25
23.77 8.91 16.11 11.69 5.65 17.66 13.22
PA-FS-EN $\times$ 75.23 36.12 39.61 49.00 44.10 45.97 48.81
48.94 9.58 16.49 14.33 7.73 19.05 19.42
Table 3: The attack success rate (ASR) defense against PA-EN and PA-FS-EN on the SST-2 dataset with LLAMA-2.
Defenses PA-EN PA-FS-EN
Vanilla 66.77 48.39
LLAMOS 21.18 37.82
Table 4: Standard accuracy and average robust accuracy on LLAMA-2 defense against three types of PromptAttack: Character-level, Word-level, and Sentence-level attacks.
FS SA RA Character Word Sentence
$\times$ 92.18 47.93 83.59 85.03 84.62
92.18 30.42 85.16 79.13 68.96
Table 5: Standard accuracy and robust accuracy of GPT-3.5 against PromptAttack. The first three rows report average accuracy across all attacks; the remaining rows give robust accuracy for each individual attack.
Methods FS SST-2 RTE QQP QNLI MNLI-mm MNLI-m Avg.
Standard - 97.66 80.47 75.78 66.41 66.41 71.87 76.43
without $\times$ 42.97 52.93 47.66 39.65 37.50 40.23 43.49
LLAMOS 24.22 51.37 45.70 33.79 37.11 38.87 38.51
LLAMOS $\times$ 82.52 77.17 67.93 61.95 67.44 63.15 70.03
72.38 76.99 66.72 60.68 65.27 62.62 67.44
C1 $\times$ 96.09 81.25 72.66 63.28 69.53 68.75 75.26
96.88 81.25 66.41 64.84 68.75 66.41 74.09
C2 $\times$ 75.78 74.22 67.19 64.84 68.75 61.72 68.75
83.59 77.34 63.28 60.16 67.97 58.59 68.49
C3 $\times$ 89.06 82.81 71.48 62.50 74.22 63.28 73.89
66.41 81.25 73.44 64.84 70.31 69.53 70.96
W1 $\times$ 85.01 79.66 70.50 61.25 68.67 65.79 71.81
80.18 77.93 69.92 59.19 64.63 63.40 69.21
W2 $\times$ 81.37 77.81 67.86 63.42 67.83 64.87 70.53
80.04 75.77 66.15 61.18 64.28 63.31 68.46
W3 $\times$ 75.79 75.12 69.57 63.76 64.85 60.91 68.33
64.48 75.00 68.17 61.69 63.00 60.51 65.47
S1 $\times$ 74.45 75.48 63.86 58.65 62.66 59.18 65.71
71.18 72.76 64.01 56.89 61.79 58.18 64.14
S2 $\times$ 85.83 74.83 64.68 59.90 65.33 62.70 68.88
58.77 76.53 64.52 59.01 63.39 61.08 63.88
S3 $\times$ 79.31 73.30 63.57 59.96 65.10 61.12 67.06
49.86 75.06 64.59 58.31 63.25 62.61 62.28
5.2 Results
Performance on ASR: We evaluate the ASR against PromptAttack-EN and PromptAttack-FS-EN on the GLUE datasets with GPT-3.5 [22]. As shown in Table 2, our method significantly reduces the ASR of both PromptAttack-EN and PromptAttack-FS-EN across all tasks. Specifically, our method achieves an average ASR reduction of 29.33% and 29.39%, respectively. These results demonstrate that LLAMOS is effective in defending against adversarial textual attacks. Additionally, we also evaluate the performance of LLAMOS on the SST-2 dataset with LLAMA-2 [32], as shown in Table 3. The results are similar to the previous experiments. The ASR of PromptAttack-EN and PromptAttack-FS-EN is significantly reduced by 45.59% and 10.57%, respectively.
Performance on RA: We evaluate the RA on the SST-2 dataset with LLAMA-2 against three types of PromptAttack: character, word, and sentence attacks. In Table 4, the first two columns represent the standard accuracy (SA) and robust accuracy (RA) without defense, while the last three columns represent the robust accuracy with LLAMOS. Under strong attacks, the accuracy of the target LLM decreased from 92.18% to 30.42%. LLAMOS can effectively defend against adversarial textual attacks, significantly improving the robust accuracy. Specifically, the lowest robust accuracy reaches 86.96%. Additionally, we conduct comprehensive experiments across nine types of attacks and six tasks with GPT-3.5, as shown in Table 5. LLAMOS can effectively defend against character-level attacks, achieving results on C1 and C3 that closely match the standard accuracy.
Table 6: Comparison on robust accuracy against C3 attack with ICL.
ICL C3 C3 with FS
$\times$ 89.06 66.41
97.66 92.19
Table 7: Robust accuracy of agents over successive rounds (R.) in the adversarial system. Each column shows accuracy after each attack–defense iteration.
Iterations R. 1 R. 2 R. 3 R. 4 R. 5
Defense 96.09 84.38 83.59 91.41 90.63
Attack 56.25 43.53 45.93 50.34 28.13
Table 8: Comparison between clean, adversarial, and purified examples. When incorrectly classified, the modified area is highlighted with a yellow background.
Method [Input] Prediction
Target LLM Clean (Original): a fast , funny , highly enjoyable movie . Positive
After Attack Adversarial: a fast , funny , highly enjoyable movie . :( Negative
LLAMOS Purified: a fast , funny , highly enjoyable movie . :( Negative
+ICL Purified: a fast , funny , highly enjoyable movie .. Positive
The purified and adversarial examples generated in the adversarial system.
Round 1 Purified: You don’t have to know music to appreciate the film’s easygoing blend of tragedy and romance. Positive
Adversarial: Youdo n’thave to know music to appreciate … Negative
Round 2 Purified: You don’t have to know music to appreciate … Positive
Adversarial: You don’t have to know music todepreciate Negative
Round 3 Purified: You don’t have to know music to appreciate … Positive
Adversarial: You don’t have to know music todepreciate Negative
LLAMOS with In-Context Learning: The C3 attack [41] is a very obvious attack that adds up to two extraneous characters to the end of the sentence, as shown in Table 8. However, our method only achieves robust accuracies of 89.06% for the C3 attack and 66.41% for the C3-FS attack. To further improve the robustness, we introduce ICL to enhance the performance of the defense agent. As shown in Section 5.2, the defense agent with ICL significantly improves the robust accuracy against the C3 attack, achieving a robust accuracies of 97.66% for the C3 attack and 92.19% for the C3-FS attack.
Analysis of Adversarial System: We conduct experiments with an adversarial system and evaluate the robust accuracy against adversarial examples generated by the attack agent over multiple iterations. As shown in Table 7, the defense agent initially achieves a robust accuracy of 96.09% in the first round of confrontation. However, after the purified examples are re-attacked by the attack agent, the robust accuracy decreases to 56.25%. The defense agent then purifies these adversarial examples again, leading to an increase in robust accuracy, but it will decrease once more by subsequent attacks. This continual fluctuation in robust accuracy is a common phenomenon in adversarial training [5]. Upon reviewing the generated texts, we observe that after several rounds of confrontation, both the defense agent and attack agent may generate the same sentences as previous ones, resulting in a potential infinite loop, as shown in Table 8. This is an interesting phenomenon that requires further investigation, particularly strategies to disrupt such loops.
5.3 Discussion
5.3.1 The Advantages of LLAMOS
As emphasized by the experimental results presented in the Results Section and Table 12, LLAMOS significantly enhances performance across various tasks and attacks with LLAMA-2 and GPT-3.5. Additionally, the defense agent in LLAMOS is a plug-and-play module, serving as a pre-processing step. Through ICL, the defense agent can be continuously optimized to defend against emerging attacks. This invisibly resolves a major challenge in adversarial robustness: Due to the significant differences between different attacks, the model trained on specific attacks often fails to generalize to other unseen attacks [8]. As a result, continuous fine-tuning is needed to stay robust against emerging attacks, but this incurs substantial costs [6, 3], and the emergence of attack techniques continually makes it impractical to retrain the model against attacks. In contrast, our method can effectively enhance the robustness through ICL without adjusting the parameters of the LLMs, which is undoubtedly a significant advantage.
5.3.2 The Challenges in LLM-based Defense
The defense agent aims to purify adversarial inputs, but it is difficult to distinguish between natural examples and adversarial examples in some cases. As shown in Table 12.4, the attacker altered the original meaning by inserting ‘not’, rendering the adversarial example indistinguishable from a natural example, resulting in the defense agent failing to generate the correct sentence. Although we hope that the defense agent can observe the sentences like humans, it presents a huge challenge at present. Unlike the attacker or humans, the defense agent lacks access to the original label of the input sentence. In addition, even successful purifications can be re-attacked. As shown in Table 7, the malicious LLMs can embed specific system prompts to influence the output, which is unbeknownst to users; they can add ‘:)’ to each input sentence for prediction rather than predicting the original sentence. In this case, the defense agent also fails to defend. This issue is also an important problem in traditional adversarial training [5], and no method has completely resolved this issue. Nonetheless, as previously discussed, LLMs offer advantages not available to traditional DNNs, and we have naturally solved one challenge in adversarial robustness, which is that the model can adapt to new attacks. Hence, future advancements may resolve adversarial issues of attack and defense within LLM frameworks, representing a challenging but promising research direction.
6 Conclusion
In this paper, we propose LLAMOS, a novel LLM-based defense technique designed to purify adversarial examples before feeding them into the target LLM. The defense agent within LLAMOS operates as a plug-and-play module that functions effectively as a pre-processing step without requiring retraining of the target LLM. We conduct extensive experiments across various tasks and attacks with LLAMA-2 and GPT-3.5. The results demonstrate that LLAMOS can effectively defend against adversarial attacks. Furthermore, we discuss certain existing shortcomings and challenges, which we aim to address in future research.
Appendix 0.A GLUE Dataset
0.A.0.1 Stanford Sentiment Treebank
SST-2 [28] is a single-sentence classification task, includes sentences from movie reviews and their sentiment annotations by humans. This task involves determining the sentiment of a given sentence, categorized into two types: positive and negative.
0.A.0.2 Recognizing Textual Entailment
RTE [2] is a binary classification task, where the goal is to determine whether a given sentence entails another sentence.
0.A.0.3 Quora Question Pairs
QQP [38] is a collection of question pairs from the community question-answering website Quora. The task is to determine whether a pair of questions are semantically equivalent.
0.A.0.4 Question-answering NLI
QNLI [23] is a natural language inference task, converted from another dataset, The Stanford Question Answering Dataset. The task is to determine whether a question and a sentence are entailed or not.
0.A.0.5 Multi-Genre Natural Language Inference
MNLI [39] is a collection of sentence pairs with textual entailment annotations. Given a premise sentence and a hypothesis sentence, the task is to predict whether the premise entails the hypothesis, contradicts the hypothesis, or neither. MNLI consists of texts from various domains and is divided into two versions: MNLI-m, where training and test datasets share the same sources, and MNLI-mm, where they differ.
Table 9: The task descriptions of the GLUE dataset.
Datasets [Task Description]
SST-2 Analyze the tone of this statement and respond with either ‘positive’ or ‘negative’.
RTE Are the following two sentences entailment or not_entailment? Answer me with ‘entailment’ or ‘not_entailment’, just one word.
QQP Are the following two questions equivalent or not? Answer me with ‘equivalent’ or ‘not_equivalent’.
QNLI Given the question and context provided, determine if the answer can be inferred by choosing ‘entailment’ or ‘not_entailment’.
MNLI-mm Does the relationship between the given sentences represent entailment, neutral, or contradiction? Respond with ‘entailment’, ‘neutral’, or ‘contradiction’.
MNLI-m Does the relationship between the given sentences represent entailment, neutral, or contradiction? Respond with ‘entailment’, ‘neutral’, or ‘contradiction’.
Table 10: The label list and input description (EEC = Each example contains). For instance, EEG one ‘sentence’ = Each example contains one ‘sentence’.
Datasets [Label List] [Input Description]
SST-2 [‘positive’, ‘negative’] EEC one ‘sentence’.
RTE [‘entailment’, ‘not_entailment’] EEC ‘sentence1’ and ‘sentence2’.
QQP [‘equivalent’, ‘not_equivalent’] EEC ‘question1’ and ‘question2’.
QNLI [‘entailment’, ‘not_entailment’] EEC ‘question’ and ‘sentence’.
MNLI-mm [‘entailment’, ‘neutral’, ‘contradiction’] EEC ‘premise’ and ‘hypothesis’.
MNLI-m [‘entailment’, ‘neutral’, ‘contradiction’] EEC ‘premise’ and ‘hypothesis’.
Appendix 0.B PromptAttack and More Comparasion Results
PromptAttack [41] modifies the clean examples at the character level, word level or sentence level. The specific guidance as shown in Table 11. Few-shot (FS): Provide examples that match the task description of adversarial attacks to LLMs to help the LLMs understand the task. Ensemble (EN): Utilize a collection of adversarial examples with various levels of perturbation, select the examples that are most likely to deceive the LLM as output.
Table 11: Perturbation instructions at the character, word, and sentence levels, respectively. This table is referenced from [41].
Level Abbre. #perturbation_instruction
Character C1 Choose at most two words in the sentence, and change them so that they have typos.
C2 Change at most two letters in the sentence.
C3 Add at most two extraneous characters to the end of the sentence.
Word W1 Replace at most two words in the sentence with synonyms.
W2 Choose at most two words in the sentence that do not contribute to the meaning of the sentence and delete them.
W3 Add at most two semantically neutral words to the sentence.
Sentence S1 Add a randomly generated short meaningless handle after the sentence, such as @fasuv3.
S2 Paraphrase the sentence.
S3 Change the syntactic structure of the sentence.
Table 12: Comparison between clean, adversarial, and purified examples. When incorrectly classified, the modified area is highlighted with a yellow background.
[Input Examples] Prediction
1. Clean Example: it ’s a charming and often affecting journey . Positive
Adv. Example: it ’s a charming and oftenaffecctingjourney . Negative
Purified Example: it’s a charming and often affecting journey. Positive
2. Clean Example: … the movie is just a plain old monster Negative
Adv. Example: … the movie is just a plain old monster:) Positive
Purified Example: … the movie is just a plain old monster . Negative
3. Clean Example: the year’s most intriguing explorations of alientation . Positive
Adv. Example: the year’s most intriguing explorations of alientation .:( Negative
Purified Example: the year’s most intriguing explorations of alientation.. Positive
4. Clean Example: it ’s a charming and often affecting journey . Positive
Adv. Example: it ’snota charming and often affecting journey .@fasuv3 Negative
Purified Example: it ’snota charming and often affecting journey . Negative
5. Clean Example: … to be kind of heartwarming, nonetheless. Positive
Adv. Example: … to be kind of heartwarming, nonetheless.@kjdjq2. Negative
Purified Example: … to be kind of heartwarming, nonetheless.. Positive